A ball is thrown at 4.7 m/s from the top of a 12 m cliff at an angle of 25 degrees below the horizontal. Determine the speed of the ball when it reaches the ground


1
Expert's answer
2021-03-31T09:32:29-0400

Let's first find the time that the ball takes to reach the ground:


"y=y_0-v_0tsin\theta-\dfrac{1}{2}gt^2,""0=12-4.7t\cdot\ sin25^{\circ}-4.9t^2,""4.9t^2+1.98t-12=0."

This quadratic equation has two roots: "t_1=1.37\ s" and "t_2=-1.78\ s". Since time can't be negative the correct answer is "t=1.37\ s".

The horizontal component of the ball's velocity doesn't change during the flight and can be found as follows:


"v_x=v_0cos\theta=4.7\ \dfrac{m}{s}\cdot cos25^{\circ}=4.26\ \dfrac{m}{s}."

The vertical component of the ball's velocity can be found as follows:


"v_y=v_0sin\theta-gt,""v_y=4.7\ \dfrac{m}{s}\cdot sin25^{\circ}-9.8\ \dfrac{m}{s^2}\cdot1.37\ s=-11.44\ \dfrac{m}{s}."

Finally, we can find the speed of the ball when it reaches the ground from the Pythagorean theorem:


"v=\sqrt{v_x^2+v_y^2}=\sqrt{(4.26\ \dfrac{m}{s})^2+(-11.44\ \dfrac{m}{s})^2}=12.2\ \dfrac{m}{s}."